OpenStudy (anonymous):

hello. everyone, Please answer for me. Write as a fraction. 30:48

OpenStudy (anonymous):

5/8

OpenStudy (anonymous):

do u mean 30/48 if yes then its 5/8

OpenStudy (anonymous):

30:48 = 30/48 = 10/16 = 5/8

OpenStudy (anonymous):

helpful? :)

OpenStudy (anonymous):

yes. thank you very much

OpenStudy (anonymous):

your welcome :)